A train ticket has a perimeter of 34 centimeters. Its area is 70 square centimeters. What are the dimensions of the ticket?

Answer by ewatrrr(24785)

 
Hi,

P =2L +2w = 34
L + w = 17
L = (17-w)
A = Lw = (17-w)w = 70
w^2 - 17w + 70 = 0
(w-10)(w -7) = 0
Dimensions: 7cm by 10cm